# Anomaly of Subsystem Symmetries in Exotic and Foliated $BF$ Theories

### v1

@@ 4 Change of Foliation Structure
Here we change the foliation structure from two foliations $e^{k}=dx^{k}\,(k=1,2)$ to three foliations $e^{k}=dx^{k}\,(k=1,2,3)$. By adding gauge fields with $k=3$ and modifying the gauge transformations, we can easily construct the 3+1d foliated SSPT phase with three foliations. The SSPT phase with a boundary also cancels the mixed ’t Hooft anomaly of the 2+1d foliated $BF$ theory on the boundary. The situation where other (exotic) SSPT phases with different foliations cancel the same anomaly appears in [25].
@@ 4.1 Foliated SSPT phase with three foliations
where $\hat{u}^{k}$ is a $x^{k}$-dependent gauge parameter valued in integers, $u^{k}dx^{k}$ is a (0+1)-form gauge parameter, $s$ is a zero-form gauge parameter, and $\hat{v}$ is the gauge parameter of $p$. Then, the 3+1d foliated SSPT phase with three foliations is written as
@@
In the case of the subsystem symmetries, the 3+1d SSPT phases with two foliations and three foliations represent the same phases from an anomaly point of view. The rotational symmetry of the SSPT phase with two foliations is the 90 degree rotation $\mathbb{Z}_{4}$ with respect to $(x^{1},x^{2})$ and the $x^{3}$ direction has no foliation. On the other hand, the rotational symmetry of the SSPT phase with three foliations is the 90 degree rotation $S_{4}$ with respect to $(x^{1},x^{2},x^{3})$. On the boundary $x^{3}=0$, both reproduce the 90 degree rotational symmetry $\mathbb{Z}_{4}$ in the exotic/foliated $BF$ theory in 2+1 dimensions. Furthermore, the SSPT phase with three foliations (4.13) is smoothly deformed into the one with two foliations (4.1) while preserving the rotational symmetry $\mathbb{Z}_{4}$ under the deformation $e^{3}=dx^{3}\rightarrow 0$, that is $\hat{\beta}^{3}$, $\beta^{3}\wedge dx^{3}$, $C^{3}\wedge dx^{3}$ and $\hat{C}^{3}$ go to 00.131313Precisely, we must consider the version where the bulk gauge fields $\hat{c}$ are not dynamical. This fact implies that we cannot apply the same discussion cannot be applied to the exotic SSPT phases. Therefore, an ’t Hooft anomaly of a subsystem symmetry in a fractonic QFT is considered to correspond to a certain deformation class of the SSPT phases and foliation structures. This gives an implication for the characterization of ’t Hooft anomalies of subsystem symmetry.
@@ 4.2 Exotic SSPT phase with three foliations
where $\hat{s}^{k}$ is a $x^{k}$-dependent gauge parameter valued in integers, and $s$ is a gauge parameter in $\bm{1}$, and we have
@@ 5 Conclusion
In this work, we have discussed the mixed ’t Hooft anomaly of subsystem symmetry in the exotic and foliated $BF$ theories in 2+1 dimensions and the SSPT phases in 3+1 dimensions that cancel it via the anomaly inflow. We have constructed the exotic and foliated SSPT phases with two and three foliations respectively by using the foliated-exotic duality. Along the way, we have shown the non-topological operator can be considered as a symmetry-like operator. We have also seen that both of the SSPT phase with two foliations and three foliations match the same ’t Hooft anomaly of the exotic/foliated $BF$ theory, and have pointed out that this fact may be a clue for characterizing ’t Hooft anomalies of subsystem symmetry.

### v2

@@ 4 Change of Foliation Structure
Here we change the foliation structure from two foliations $e^{k}=dx^{k}\,(k=1,2)$ to three foliations $e^{k}=dx^{k}\,(k=1,2,3)$. By adding gauge fields with $k=3$ and modifying the gauge transformations, we can easily construct the 3+1d foliated SSPT phase with three foliations. The SSPT phase with a boundary also cancels the mixed ’t Hooft anomaly of the 2+1d foliated $BF$ theory on the boundary. The situation where other (exotic) SSPT phases with different foliations cancel the same anomaly appears in [25, 30].
@@ 4.1 Foliated SSPT phase with three foliations
where $\hat{u}^{k}$ is a $x^{k}$-dependent integer-valued gauge parameter, $u^{k}dx^{k}$ is a (0+1)-form gauge parameter, $s$ is a zero-form gauge parameter, and $\hat{v}$ is the gauge parameter of $p$. Then, the 3+1d foliated SSPT phase with three foliations is written as
@@
In the case of the subsystem symmetries, the 3+1d SSPT phases with two foliations and three foliations represent the same phases from an anomaly point of view. The rotational symmetry of the SSPT phase with two foliations is the 90 degree rotation $\mathbb{Z}_{4}$ with respect to $(x^{1},x^{2})$ and the $x^{3}$ direction has no foliation. On the other hand, the rotational symmetry of the SSPT phase with three foliations is the 90 degree rotation $S_{4}$ with respect to $(x^{1},x^{2},x^{3})$. On the boundary $x^{3}=0$, both reproduce the 90 degree rotational symmetry $\mathbb{Z}_{4}$ in the exotic/foliated $BF$ theory in 2+1 dimensions. Furthermore, the SSPT phase with three foliations (4.13) is smoothly deformed into the one with two foliations (4.1) while preserving the rotational symmetry $\mathbb{Z}_{4}$ under the deformation $e^{3}=dx^{3}\rightarrow 0$, that is $\hat{\beta}^{3}$, $\beta^{3}\wedge dx^{3}$, $C^{3}\wedge dx^{3}$ and $\hat{C}^{3}$ go to 00.141414Precisely, we must consider the version where the bulk gauge fields $\hat{c}$ are not dynamical. This fact implies that we cannot apply the same discussion to the exotic SSPT phases. Therefore, an ’t Hooft anomaly of a subsystem symmetry in a fractonic QFT is considered to correspond to a certain deformation class of the SSPT phases and foliation structures. This gives an implication for the characterization of ’t Hooft anomalies of subsystem symmetry.
@@ 4.2 Exotic SSPT phase with three foliations
where $\hat{s}^{k}$ is a $x^{k}$-dependent integer-valued gauge parameter, and $s$ is a gauge parameter in $\bm{1}$, and we have
@@ 5 Conclusion
In this work, we have discussed the mixed ’t Hooft anomaly of subsystem symmetry in the exotic and foliated $BF$ theories in 2+1 dimensions and the SSPT phases in 3+1 dimensions that cancel it via the anomaly inflow. We have constructed the exotic and foliated SSPT phases with two and three foliations respectively by using the foliated-exotic duality. Along the way, we have shown the non-topological defect that describes a fracton can be considered as a symmetry-like operator. We have also seen that both of the SSPT phase with two foliations and three foliations match the same ’t Hooft anomaly of the exotic/foliated $BF$ theory, and have pointed out that this fact may be a clue for characterizing ’t Hooft anomalies of subsystem symmetry.